# Quantum nonlinear effects in the number-conserving analogue gravity of Bose-Einstein condensates

### v1

@@ III.1 Linearized fluctuations around a background
Here, the scaling of the fields with the number of particles is as follows,
@@
and so on. For later convenience, we can establish a one-to-one correspondence with the subscript of each component of the density perturbation with the relevant scaling with $N^{1/2}$. For example, the subscript $0\rightarrow\mathcal{O}(N^{1/2})$ and similarly for other terms in descending order. Note that $f_{1}=\mathcal{O}(N^{0})$ and $f_{3}=\mathcal{O}(N^{-1})$ defined earlier in (LABEL:f1f2def) follow the same scaling convention, i.e. an index $\alpha$ implies of ${\cal O}(N^{(1-\alpha)/2})$.
At this point, if we choose to follow the Unruh paradigm [4, 2], then it would suffice to keep only ${\mathcal{O}}(\epsilon)$ terms, and we can study the dynamics of the perturbed phase $\delta\theta$ around the background $\rho_{0},\theta_{0}$, using that the continuity equation remains satisfied. Then it is possible to describe the dynamics of the (first-order) fluctuation field $\delta\theta_{1}$ as a massless Klein-Gordon equation in an effective curved spacetime
@@
The components of the metric $g_{\mu\nu}$ depend on the background variables $\rho_{0},\theta_{0}$, where the speed of sound in this background is $c^{2}_{s0}=\frac{U}{\rho_{0}}{m}$.
@@ III.2 Second-order fluctuations around a background
Let us first assume expansions of the form of Eqs. (22) and (23)) are possible, up to the indicated quadratic order. Then from the modified continuity equation (19), and the Euler equation 20, we obtain the following set of equations:
$\partial_{t}(\rho_{0}+\epsilon\delta\rho_{1})+\nabla\cdot\Big{(}(\rho_{0}+\epsilon\delta\rho_{1})(\frac{\nabla\theta_{0}}{m}+\epsilon\frac{\nabla\delta\theta_{1}}{m})\Big{)}=f_{1}(\rho_{0},\theta_{0})+\epsilon\frac{\partial f_{1}}{\partial\rho_{c}}|_{\rho_{0}}\delta\rho_{1}+\epsilon\frac{\partial f_{1}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1},$ (26)
$\partial_{t}(\theta_{0}+\epsilon\delta\theta_{1})+U(\rho_{0}+\epsilon\delta\rho_{1})+\frac{1}{2m}(\nabla(\theta_{0}+\epsilon\delta\theta_{1}))^{2}+V(x)$
$+f_{3}(\rho_{0},\theta_{0})+\epsilon\frac{\partial f_{3}}{\partial\rho_{c}}|_{\rho_{0}}\delta\rho_{1}+\epsilon\frac{\partial f_{3}}{\partial\theta_{c}}|_{\theta_{0}}\delta\theta_{1}=0.$ (27)
As we will argue later in detail, the two above equations define the new background on which we will consider the propagation of nonlinear fluctuations. Then the new equations governing the perturbations are
$\epsilon^{2}\partial_{t}\delta\rho_{2}+\nabla\cdot\Big{(}\epsilon^{2}\delta\rho_{2}(\frac{\nabla\theta_{0}}{m}+\epsilon\frac{\nabla\delta\theta_{1}}{m})+\epsilon^{2}(\rho_{0}+\epsilon\delta\rho_{1})\frac{\nabla\delta\theta_{2}}{m}\Big{)}=\epsilon^{2}\frac{\partial^{2}f_{1}}{\partial\rho_{c}^{2}}|_{\rho_{0}}\delta\rho_{2}+\epsilon^{2}\frac{\partial^{2}f_{1}}{\partial\theta_{c}^{2}}|_{\theta_{0}}\delta\theta_{2},$ (28)
$\epsilon^{2}\partial_{t}\delta\theta_{2}+\epsilon^{2}U\delta\rho_{2}+\epsilon^{2}\frac{1}{m}\nabla(\theta_{0}+\epsilon\delta\theta_{1})\cdot\nabla\delta\theta_{2}+\epsilon^{2}\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}|_{\rho_{0}}\delta\rho_{2}+\epsilon^{2}\frac{\partial^{2}f_{3}}{\partial\theta_{c}^{2}}|_{\theta_{0}}\delta\theta_{2}=0.$ (29)
@@
We now consider the following two combinations of fluid variables,
@@
where we truncated in the second line the series in Eqs. (22) and (23) after quadratic order. We assume that these combinations satisfy the dynamical fluid Eqs. (19) and (20)) for two different boundary conditions, such that the expansions above are convergent, which in turn requires $\epsilon^{2}\delta\rho_{2}\ll\epsilon\delta\rho_{1}\ll\rho_{0}$. This also implies that we are not separately demanding $\rho_{0},\theta_{0}$ to follow any dynamical equations, at least formally. Then the condition on $\rho_{0},\theta_{0}$ is simply that they are two sufficiently slowly varying “background” functions, which are not necessarily solutions to the fluid dynamical equations, but can be chosen, for example, according to the ease of interpretation of experiments. Namely, such that the first and second order perturbative corrections on top of the background satisfy the fluid-dynamical equations.
We can consider of course also fluctuations higher than quadratic order and, in a similar way, demand that we can suitably define a background for the next-order density and phase fluctuation. In a minimal self-consistent analogue gravity model, as a consequence of the backreaction due to enforcing number conservation, the perturbed fluid variables have however to be at least of quadratic order, i.e., should scale here as $\mathcal{O}(N^{-1/2})$, to be properly treated as a perturbation.
@@
$\begin{split}\partial_{t}\Big{(}\frac{1}{\tilde{U}}\Big{[}\partial_{t}\delta\theta_{2}+v_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}\Big{)}+\nabla\cdot\Bigg{(}\frac{1}{\tilde{U}}v_{b}\Big{[}\partial_{t}\delta\theta_{2}+v_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}-\frac{\rho_{b}}{m}\nabla\delta\theta_{2}\Bigg{)}\\ -\frac{\ddot{f}_{1}}{\tilde{U}}\Big{(}\partial_{t}\delta\theta_{2}+v_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{)}+f^{\prime\prime}_{1}\delta\theta_{2}=0.\end{split}$ (32)
@@
$v_{b}=\frac{\nabla(\theta_{0}+\epsilon\delta\theta_{1})}{m},$ (33)
@@
$\tilde{U}=U+\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}|_{\rho_{0}}.$ (34)
@@ III.3 Lorentz-invariant wave equation for second-order perturbations
Here, we have defined the modified derivative operator as (see for the definition of $\mathcal{A}_{\mu}$ Eq. (40) below)
@@
$\mathbf{g}_{tt}=-\sqrt{\frac{\rho_{b}}{m\tilde{U}}}\left(\frac{\tilde{U}\rho_{b}}{m}-v_{b}^{2}\right),$ (37)
$\mathbf{g}_{ti}=-\sqrt{\frac{\rho_{b}}{m\tilde{U}}}(v_{b})_{i},\qquad\mathbf{g}_{ij}=\sqrt{\frac{\rho_{b}}{m\tilde{U}}}\delta_{ij},$ (38)
@@
$\sqrt{-\mathbf{g}}=\sqrt{\frac{\rho_{b}^{3}}{m^{3}\tilde{U}}}.$ (39)
Here, the index $i$ represents the spatial coordinates. We have defined the components of a vector field $\mathcal{A}^{\mu}$ as
$\mathcal{A}^{\mu}=\Big{(}\frac{f^{\prime\prime}_{3}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{f^{\prime\prime}_{3}v_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)},$ (40)
and the components of the field $\mathcal{B}^{\mu}$ as
$\mathcal{B}^{\mu}=\Big{(}\frac{\ddot{f}_{1}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{\ddot{f}_{1}v_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)}.$ (41)
@@
In summary, the dynamics of the nonlinear phase fluctuations can be written, to second order in the perturbations, as a generalized massive Klein-Gordon equation in a curved spacetime, coupled with two emergent vector fields, all with respect to the background $\rho_{b},v_{b}$.
@@
In our context of the BEC-based analogue gravity model incorporating number conservation, the above corrections to the Unruh paradigm can be readily understood by considering a system where the interaction is turned on suddenly in an initially noninteracting gas [35], also see the Appendix. Then, at sufficiently early time, when the number of depleted particles is very small, the correction due to backreaction is also negligible, and the linear-regime Unruh paradigm metric is an excellent description. However, as the system evolves, the number of depleted particles is no longer negligible, and the corrected metric will afford a better description.
@@
The equation (35) is not invariant under time reversal $t\rightarrow-t$. The presence of a first-order derivative in time breaks the time reversal symmetry, which in turn can be traced back to the existence of particle exchange between the condensed and the noncondensed parts of the system, which is fully respected in our construction of a number-conserving theory. The manifestation of the breaking of time reversal is an imaginary quasiparticle frequency, cf. Eq. (45) below.
III.4 Comparison to classical nonlinear approach
@@
At this point, it is appropriate to draw a comparison between the formulation we have adopted here and that proposed in the classical analysis of [40], where the propagation of a general nonlinear wave in a fluid medium described by the coupled continuity and Euler equations was considered from the viewpoint of analogue gravity.
@@
It was shown that the basic premise of analogue gravity, i.e., the propagation of linear perturbations on a slowly varying background, can be described by an effective metric. This description however breaks down after a certain time scale due to the inherent nonlinear nature of the fluid motion. It is nevertheless possible to obtain a spacetime metric description provided we take the backreaction of the nonlinear motion into account and consequently modify the background itself. In the present paper, we have taken a weaker restriction on the nature of nonlinearity in the fluid under consideration. To be more specific, distinct from [40], we have assumed that the perturbation expansion in (22) is convergent, i.e., the successive terms in this expansion are gradually of sufficiently lower order than the previous one, an assumption that was not a priori taken in [40]. Also our analysis will break down after a certain timescale at which the fully nonperturbative nature of the wave propagation can no longer be neglected. The exact relation between this timescale, and the failure of the Bogoliubov approximation, i.e., when the number of condensate particles becomes of order the number of noncondensed particles, should be an interesting topic to pursue in the future.
@@
We have summarized the features of the various approaches to analogue gravity which have been taken since its inauguration in table1, together with a few selected references. The combination of generality and degree of control possible within our approach (which can be extended to higher orders of powers of $N^{-1/2}$) becomes apparent from this comparison.
Approach Background considered Number-conserving References
Leading-order GPE (5)    Linearized on top of $(\rho_{0},v_{0})$ No [8, 9, 12]
Number-conserving ansatz (3) Full fluid density and velocity [^1] In Refs. [29],and [35] the importance of full number conservation was pointed out. However, no consequences for the analogue metric      to any given order in the large $N$ expansion were considered. Yes [29, 35]
Nonlinear waves in classical fluid Nonperturbative fluctuation (redefined background) [^2] The formulation in [40] is for a generic barotropic, non-viscous, and irrotational fluid, with a BEC as a particular case. However, the redefined background in [40] is not the $(\rho_{b},v_{b})$ employed here. N/A [40]
Corrected GPE (9) Second-order on background $(\rho_{b},v_{b})$ Yes This work
Table 1: A table comparing formulations used to construct analogue gravity models in quantum and classical fluids.

### v2

@@ III.1 Linearized fluctuations around a background
Here, the scaling of the density contributions with the number of particles is as follows,
@@
and so on. For later convenience, we can thus establish a one-to-one correspondence with the subscript of each component of the density perturbation with the relevant scaling with $N^{1/2}$. For example, the subscript $0\rightarrow\mathcal{O}(N^{1/2})$ and similarly for other terms in descending order, i.e., an index $\alpha$ implies being of ${\cal O}(N^{(1-\alpha)/2})$. Note also that $f_{1}=\mathcal{O}(N^{0})$ and $f_{3}=\mathcal{O}(N^{-1})$ defined earlier in (LABEL:f1f2def) follow the same scaling convention, but that the phase perturbation index in (25) does not indicate a scaling with powers of $N^{1/2}$.
We stress here for clarity that we are dealing with two expansion parameters in terms of small quantities, where $N^{-1/2}$ is the mean-field expansion parameter and $\epsilon$ is the parameter indicating the amplitude perturbation order of the fields. If we choose to follow the Unruh paradigm [4, 2], then it would suffice to keep only ${\mathcal{O}}(\epsilon)$ terms, and we can study the dynamics of the perturbed phase $\delta\theta$ around the background $\rho_{0},\theta_{0}$, using that the continuity equation remains satisfied. Then it is possible to describe the dynamics of the (first-order) fluctuation field $\delta\theta_{1}$ as a massless Klein-Gordon equation in an effective curved spacetime
@@
The components of the metric $g_{\mu\nu}$ depend on the background variables $\rho_{0},\theta_{0}$, where the speed of sound in this background is $c^{2}_{s0}=U\rho_{0}$.
@@ III.2 Second-order fluctuations around a background
Let us first assume perturbation amplitude expansions as in Eqs. (24) and (25)) are possible, up to the indicated quadratic order. Then from the modified continuity equation (21), and the Euler equation 22, we obtain the following set of equations:
$\partial_{t}(\rho_{0}+\epsilon\delta\rho_{1})+\nabla\cdot\Big{(}(\rho_{0}+\epsilon\delta\rho_{1})(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})\Big{)}=f_{1}(\rho_{0},\theta_{0})+\epsilon\left.\frac{\partial f_{1}}{\partial\rho_{c}}\right|_{0}\delta\rho_{1}+\epsilon\left.\frac{\partial f_{1}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}+\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\delta\rho_{1}\delta\theta_{1},$
$\partial_{t}(\theta_{0}+\epsilon\delta\theta_{1})+U(\rho_{0}+\epsilon\delta\rho_{1})+\frac{1}{2}(\nabla(\theta_{0}+\epsilon\delta\theta_{1}))^{2}+V+f_{3}(\rho_{0},\theta_{0})+\epsilon\left.\frac{\partial f_{3}}{\partial\rho_{c}}\right|_{0}\delta\rho_{1}+\epsilon\left.\frac{\partial f_{3}}{\partial\theta_{c}}\right|_{0}\delta\theta_{1}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}\partial\theta_{c}}\right|_{0}\delta\rho_{1}\delta\theta_{1}=0.$
where we used the abbreviation $\left.\cdots\right|_{0}\coloneqq\left.\cdots\right|_{\rho_{0},\theta_{0}}$ here and in what follows. As we will argue later in detail, the two above equations define the new background on which we will consider the propagation of nonlinear fluctuations. Then the new equations governing the perturbations are
$\epsilon^{2}\partial_{t}\delta\rho_{2}+\nabla\cdot\Big{(}\epsilon^{2}\delta\rho_{2}(\nabla\theta_{0}+\epsilon\nabla\delta\theta_{1})+\epsilon^{2}(\rho_{0}+\epsilon\delta\rho_{1})\nabla\delta\theta_{2}\Big{)}=\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\rho_{c}^{2}}\right|_{0}\delta\rho_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{1}}{\partial\theta_{c}^{2}}\right|_{0}\delta\theta_{2},$ (30)
$\epsilon^{2}\partial_{t}\delta\theta_{2}+\epsilon^{2}U\delta\rho_{2}+\epsilon^{2}\nabla(\theta_{0}+\epsilon\delta\theta_{1})\cdot\nabla\delta\theta_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}\right|_{0}\delta\rho_{2}+\epsilon^{2}\left.\frac{\partial^{2}f_{3}}{\partial\theta_{c}^{2}}\right|_{0}\delta\theta_{2}=0.$ (31)
@@
Let us consider the following combinations of variables,
@@
where we truncated in the second line the series in Eqs. (24) and (25) after second order in $\epsilon$. We assume that these combinations satisfy the dynamical fluid Eqs. (21) and (22)) for two different boundary conditions, such that the expansions above are convergent, which in turn requires $\epsilon^{2}\delta\rho_{2}\ll\epsilon\delta\rho_{1}\ll\rho_{0}$. This also implies that we are not separately demanding $\rho_{0},\theta_{0}$ to follow any dynamical equations, at least formally. Then the condition on $\rho_{0},\theta_{0}$ is simply that they are two sufficiently slowly varying “background” functions, which are not necessarily solutions to the fluid dynamical equations, but can be chosen, for example, according to the ease of interpretation of experiments. Namely, such that the first and second order perturbative corrections on top of the background satisfy the fluid-dynamical equations.
We can consider of course also fluctuations higher than of quadratic order and, in a similar fashion, demand that we can suitably define a background for the next-order density and phase fluctuations. In a minimally self-consistent analogue gravity model, as a consequence of the backreaction due to enforcing number conservation, the perturbed fluid variables have however to be at least of quadratic order, i.e., should scale here as $\mathcal{O}(N^{-1/2})$, to be properly treated as a perturbation.
@@
$\partial_{t}\Big{(}\frac{1}{\tilde{U}}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}\Big{)}+\nabla\cdot\Bigg{(}\frac{1}{\tilde{U}}{\bm{v}}_{b}\Big{[}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{]}-\rho_{b}\nabla\delta\theta_{2}\Bigg{)}\\ -\frac{\ddot{f}_{1}}{\tilde{U}}\Big{(}\partial_{t}\delta\theta_{2}+{\bm{v}}_{b}\cdot\nabla\delta\theta_{2}+f^{\prime\prime}_{3}\delta\theta_{2}\Big{)}+f^{\prime\prime}_{1}\delta\theta_{2}=0.$ (34)
@@
${\bm{v}}_{b}=\nabla(\theta_{0}+\epsilon\delta\theta_{1}),$ (35)
@@
$\tilde{U}=U+\left.\frac{\partial^{2}f_{3}}{\partial\rho_{c}^{2}}\right|_{0}.$ (36)
@@ III.3 Lorentz-invariant wave equation for second-order perturbations
Here, we have defined the modified derivative operator as
@@
Note that $A_{\mu}$ does not assume the role of a gauge field here and is defined such that it is real; see Eq. (42) below.
@@
$\mathbf{g}_{tt}=-\sqrt{\frac{\rho_{b}}{\tilde{U}}}\left(\tilde{U}\rho_{b}-{\bm{v}}_{b}^{2}\right),$ (39)
$\mathbf{g}_{ti}=-\sqrt{\frac{\rho_{b}}{\tilde{U}}}{\bm{v}}_{bi},\qquad\mathbf{g}_{ij}=\sqrt{\frac{\rho_{b}}{\tilde{U}}}\delta_{ij},$ (40)
@@
$\sqrt{-\mathbf{g}}=\sqrt{\frac{\rho_{b}^{3}}{\tilde{U}}}.$ (41)
Here, the index $i$ represents the spatial coordinates.
We have defined in Eqs. (37) and (38) the components of a vector field
$\mathcal{A}^{\mu}=\Big{(}\frac{f^{\prime\prime}_{3}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{f^{\prime\prime}_{3}{\bm{v}}_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)},$ (42)
and the components of a second vector field $\mathcal{B}^{\mu}$ as
$\mathcal{B}^{\mu}=\Big{(}\frac{\ddot{f}_{1}}{\sqrt{\mathbf{-g}}\tilde{U}},\frac{\ddot{f}_{1}{\bm{v}}_{b}}{\sqrt{\mathbf{-g}}\tilde{U}}\Big{)}.$ (43)
@@
In summary, the dynamics of the nonlinear phase fluctuations can be written, to second order in the perturbations, as a generalized massive Klein-Gordon equation in a curved spacetime, coupled with two emergent vector fields, all with respect to the background $\rho_{b},{\bm{v}}_{b}$.
@@
In our context of the BEC-based analogue gravity model incorporating number conservation, the above corrections to the Unruh paradigm can be readily understood by considering a system where the interaction is turned on suddenly in an initially noninteracting gas [36], also see the Appendix. Then, at sufficiently early time, when the number of depleted particles is very small, the correction due to backreaction is also negligible, and the Unruh paradigm metric is an excellent description. However, as the system evolves, the number of depleted particles is no longer negligible, and the corrected metric will afford a better description.
@@
The equation (37) is not invariant under time reversal $t\rightarrow-t$. The presence of a first-order derivative in time breaks the time reversal symmetry, which in turn can be traced back to the existence of particle exchange between the condensed and the noncondensed parts of the system, which is fully respected in our construction of a number-conserving theory. The manifestation of the breaking of time reversal is a complex quasiparticle frequency, cf. Eq. (47) below.
III.4 Comparison to other approaches
@@
At this point, it is appropriate to draw a comparison between the formulation we have adopted here and in particular the one proposed in the classical analysis of [41], where the propagation of a general nonlinear wave in a fluid medium described by the coupled continuity and Euler equations was considered from the viewpoint of analogue gravity.
@@
It was shown in [41] that an effective metric can be derived also when the intrinsic nonlinearity of a (classical perfect) fluid is taken into account. This description however breaks down after a certain time scale due to the inherent nonlinear nature of the fluid motion and the consequent creation of shocks. It is nevertheless possible to obtain a spacetime metric description provided we take the backreaction of the nonlinear motion into account and consequently modify the background itself. In the present paper, we have taken a weaker restriction on the nature of nonlinearity in the fluid under consideration. To be more specific, distinct from [41], we have assumed that the perturbation expansion in (24) is convergent, i.e., the successive terms in this expansion are gradually of sufficiently lower order than the previous one, an assumption that was not a priori taken in [41]. Also our analysis will break down after a certain timescale at which the fully nonperturbative nature of the wave propagation can no longer be neglected. The exact relation between this timescale, and the failure of the Bogoliubov approximation, i.e., when the number of condensate particles becomes of order the number of noncondensed particles, should be an interesting topic to pursue in the future.
Approach Background considered Number-conserving References
Leading-order GPE (5)    Linearized on top of $(\rho_{0},{\bm{v}}_{0})$ No [8, 9, 12]
Number-conserving ansatz (3) Full fluid density and velocity [^3] In Refs. [30],and [36] the importance of full number conservation was pointed out. However, no consequences for the analogue metric      to any given order in the large $N$ expansion were considered. Yes [30]
Nonlinear waves in classical fluid Nonperturbative fluctuation (redefined background) [^4] The formulation in [41] is for a generic barotropic, non-viscous, and irrotational fluid, with a BEC as a particular case. However, the redefined background in [41] is not the $(\rho_{b},{\bm{v}}_{b})$ employed here. N/A [41]
Corrected GPE (11) Second-order on background $(\rho_{b},{\bm{v}}_{b})$ Yes [36], This work
Table 1: A table comparing formulations used to construct analogue gravity models for quantum and classical fluids.
@@
We have summarized the features of various approaches to analogue gravity in quantum and classical fluids which have been taken since its inauguration, in table 1, together with a few selected references. The combination of generality and degree of control possible within our quantum many-body approach (which can be extended to higher orders of powers of $N^{-1/2}$) becomes apparent. Most importantly, we provide the first spacetime metric description of the propagation of second-order perturbations in a quantum fluid, within a number-conserving formulation of backreaction.